### v1

# Elastic Data Transfer Optimization with Hybrid Reinforcement Learning

Abstract
Modern scientific data acquisition generates petabytes of data that must be transferred to geographically distant computing clusters. Conventional tools either rely on preconfigured sessions, which are difficult to tune for users without domain expertise, or they adaptively optimize only concurrency while ignoring other important parameters. We present LDM, an adaptive data transfer method that jointly considers multiple parameters. Our solution incorporates heuristic-based parallelism, infinite pipelining, and a deep reinforcement learning based concurrency optimizer. To make agent training practical, we introduce a lightweight network simulator that reduces training time to less than four minutes and provides a $2750\times$ speedup compared to online training. Experimental evaluation shows that LDMconsistently outperforms existing methods across diverse datasets, achieving up to $9.5\times$ higher throughput compared to state-of-the-art solutions.
I Introduction
The modern scientific landscape is experiencing a data deluge, driven by advances in instrumentation, sensing technologies, and computational power that generate datasets at an unprecedented scale. This phenomenon, often referred to as “Big Data,” is commonly characterized by three dimensions: immense volume, high velocity of acquisition, and wide variety of data types. Research infrastructures such as high-energy physics experiments, large-scale genomics initiatives, and global sensor networks now routinely produce petabytes of data. For analysis, this data often must be transferred to geographically distant High-Performance Computing (HPC) clusters [ligo, ares, combustion, astronomy, des, lsst, atlas, belle2, ameriflux, esgf, esxsnmp]. Moving data at this scale requires extremely high network bandwidth. To address this need, Internet2 has already increased its backbone to 400 Gbps [internet2400Gbps], and ongoing efforts aim to extend network capacity to the terabit scale [esnet1000Gbps].
To maximize the capabilities of modern transfer infrastructures, it is essential to carefully tune a variety of parameters, such as pipelining (transferring multiple files through a single connection) [TCP_Pipeline, farkas2002], parallelism (transferring one file in parts) [R_Lee01, R_Hacker05, R_Karrer06, R_Lu05] and concurrency (transferring multiple files at a time) [kosar04, Kosar09, R_Liu10]. These adjustments can be made entirely at the application layer without altering the underlying transfer protocols, and they play a crucial role in enhancing the overall efficiency and throughput of end-to-end data transfers. Most existing approaches to maximizing bandwidth utilization in high-speed networks rely solely on concurrency. This method can be effective for datasets with uniformly sized medium-to-large files. However, it performs poorly when the dataset contains a few very large files or a large number of small files, since it does not take parallelism or pipelining into account.
The key challenge is to design a framework that remains user-friendly, since end users often lack the domain knowledge required to manually tune parameters for optimal performance. Another important consideration is that networks are not static. Bandwidth can fluctuate due to various factors, and a fixed parameter configuration may no longer remain optimal. For this reason, most recent works have modeled the problem as an online optimization task. Hasibul et al. [drl] addressed this problem using deep reinforcement learning. Although the approach is conceptually sound, it is impractical in practice because it requires day-long online training before the agent can be deployed.
This paper presents LDM, a novel adaptive data transfer method that considers not only concurrency but also pipelining and parallelism. Parallelism is addressed using a heuristic approach based on file size, while pipelining is implemented as an infinite strategy. Concurrency is optimized through a policy-driven deep reinforcement learning (DRL) agent. To eliminate the impracticality of day-long online training [drl], we design a lightweight simulator that reduces training time to as little as four minutes.
In summary, the major contributions of this work are:
  • We propose an adaptive multi-parameter framework for maximizing bandwidth utilization during data transfer. Parallelism is determined heuristically, pipelining is implemented as infinite, and concurrency is optimized using a Proximal Policy Optimization (PPO) agent. These three parameters operate transparently, requiring no user expertise. Unlike approaches that rely solely on concurrency, our method generalizes across diverse dataset file-size distributions.
  • We introduce a lightweight network simulator that emulates data transfer dynamics and enables efficient agent training. This simulator achieves a $2750\times$ speedup compared to online training.
  • We demonstrate through extensive experiments that LDM outperforms existing methods across all dataset scenarios, achieving up to $9.5\times$ higher throughput compared to state-of-the-art solutions.
II Related Work & Motivation
Trade-off Between Performance and Overhead: Both heuristic and supervised machine learning models estimate parameter values before data transfers begin, using fixed configurations throughout the entire process. This static approach prevents adaptation to changing network conditions during long-running transfers. Additionally, these solutions lack a back-off mechanism, meaning that large parameter values may overwhelm systems in various production environments. For example, if $x$ parallel TCP connections are required to fully utilize the available bandwidth, assigning $x$ connections to every new transfer becomes inefficient when multiple concurrent transfers are initiated. In a scenario with $N$ concurrent transfers, $Nx$ connections would be created, leading to diminished throughput per connection (by a factor of $N$), but significantly increased overhead (Figure LABEL:fig:cc_impacts). A more efficient alternative would be to allocate fewer connections, say, 10% of $x$ – to each new transfer, under the assumption that overlapping transfers or other processes will compete for I/O resources. For instance, Globus Online, a widely adopted data transfer service, allows tuning of three key parameters: pipelining, parallelism, and concurrency. However, in production environments, system administrators set these values heuristically and often very conservatively to minimize I/O and network contention. While this approach helps balance overhead with resource utilization, it frequently leaves networks underutilized. As a result, static configurations present a dilemma: either prioritize utilization at the cost of high overhead, or reduce overhead and risk underutilizing the network.
Long Convergence Time of Adaptive Solutions: Static solutions become suboptimal when the host or network environments deviate from initial assumption; a common occurrence in long-running transfers due to constantly evolving network dynamics. Moreover, these solutions lack an effective back-off mechanism to manage overhead. Ideally, at the start of a transfer, a system should maximize its bandwidth usage, but as new transfers join, existing transfers should release some connections. Similarly, when a transfer finishes, remaining transfers should immediately capture the freed bandwidth. However, to be useful, these solutions must quickly converge to optimal configurations at the beginning of a transfer or whenever network conditions change, minimizing the search time spent in suboptimal states. The convergence speed is directly related to the search space: with N parameters and M possible ranges per parameter, the search space is $M^{N}$. This exponential growth in the search space with each additional parameter increases convergence time significantly.
Stochastic approximation algorithms [rao2016experimental, yun2017data, esma-tcc, liu2018toward] offer promising alternatives to static solutions as they can discover optimal parameter configuration in the runtime. However, Existing solutions in this domain fall short to offer practical option due to long convergence time and failure to provide fairness and stability guarantees in shared environments. for example, Yun et al. proposed ProbData [yun2017data] to tune the number of parallel streams and buffer size for memory-to-memory TCP transfers using stochastic approximation. ProbData is able to explore the near-optimal configurations through sample transfers but it takes several hours to converge which makes it impractical to use as majority of the transfers in HPNs last less than few hours [liu2018cross]. Also, we have observed that background traffic changes drastically over several hours in shared networks [arslan2018high], so it may even fail to converge due to large variations in sample transfers.
Simplified Architecture Overwhelm Systems: The most recent studies also have implemented adaptive approaches using online black-box optimization algorithms such as Bayesian methods, Gradient Descent, and Deep Reinforcement Learning [arifuzzaman2023falcontpds, marlin, jamil2023drl]. These methods offer high utilization with reduced overhead by periodically collecting socket statistics during transfers. The black-box design allows for instant deployment across various networks, as they don’t rely on historical data. These studies finds out that, Concurrency is the most impactful parameter for data transfers. As a result, these solutions narrow the problem to optimizing the number of Concurrency, which can offset suboptimal settings in other parameters. For instance, on a 100 Gbps link, an optimally tuned connection might achieve 10 Gbps, so the solution would require 10 parallel connections. But if TCP buffer size value is suboptimal, throughput per connection could drop to 1 Gbps, prompting the optimizer to increase concurrency to 100. The optimizer can very quickly, often as little in 30 seconds, find this target value due to reduced search space. While these solutions now can easily achieved high network utilization, the increasing concurrency values might significantly degrades the performance of the existing systems and I/O processes as demonstrated in Figure LABEL:fig:cc_impacts. Although extending these solutions to multi-parameter optimization is fairly straightforward, they often perform much worse because the optimizer spends significant time searching for the optimal configuration due to exponentially increasing search space.
III LDM: Systems Design
We introduce LDM to address the throughput maximization problem in high-speed data transfer. It combines both high-level and low-level design strategies. At the high level, we introduce a novel pipeline that determines optimal, stable, and fast-converging transfer parameters. This control strategy is developed and refined using an offline simulator, which allows for efficient agent training. At the low level, we focus on the implementation of this strategy, incorporating robust mechanisms for handling pipelining and parallelism within the production system.
Figure 1: LDM optimizes concurrency, parallelism, and pipelining without being tuned by the end user.
III-A Handling Transfer Parameters in LDM
In LDM, we address the three key transfer parameters—parallelism, pipelining, and concurrency—in unique ways. The overall design is illustrated in the flowchart shown in Figure 1.
Parallelism. The conventional approach fixes the number of parallel streams per file. However, this method is problematic because it assigns the same number of streams to both a 10 KB file and a 10 GB file. Since parallelism involves splitting and merging files, unnecessary parallelism can create bottlenecks. To overcome this, we implemented a maximum chunk size–based strategy. A maximum chunk size is defined in advance, and each file is transmitted with $\lceil filesize/chunksize\rceil$ parallel streams. This ensures that large files are allocated more streams, while small files typically use only one stream, avoiding unnecessary system overhead.
Pipelining. Pipelining determines how frequently connections are closed and re-established. In LDM, we use infinite pipelining: a connection remains open until the worker is stalled by the concurrency controller or the transfer is completed. This approach allows the congestion window to grow to its full potential, avoiding the penalty of restarting from the initial window each time a connection is closed.
Concurrency. Concurrency remains the central and most critical parameter, as both pipelining and parallelism are partially dependent on it. For example, parallelism becomes ineffective when concurrency is fixed at one, and pipelining suffers if concurrency is unstable and frequently turns connections on and off. In LDM, concurrency is dynamically determined by a PPO agent, which we describe in detail in the following subsection.
III-B Utility Function
Since we model concurrency as an optimization problem, we require a utility function that captures the essential factors to guide concurrency decisions. For this purpose, we use throughput and packet loss rate as the key parameters. These are collected during the data transfer session and used to assign a score to the current concurrency level. The utility score increases only when throughput improves meaningfully. Small gains are discouraged, as increasing concurrency for negligible improvement unnecessarily burdens the CPU. For example, suppose concurrency doubles from 8 to 16 but throughput rises only from 20.0 Gbps to 20.5 Gbps. In that case, the utility score penalizes the extra streams since the marginal benefit does not justify the additional resource cost. In essence, the objective is to maximize throughput while minimizing the number of concurrent streams. The utility function is defined as:
$U(thrpt,cc)$ $=\frac{thrpt}{K^{cc}}-B\cdot plr$
Here, thrptthrpt denotes throughput, $cc$ is concurrency, and plrplr is the packet loss rate, all of which are probed during an ongoing transfer. Throughput contributes positively to the score, while the exponential penalty term $K^{cc}$ discourages excessive concurrency. The constant $K$ is a tunable parameter that balances throughput gains against resource consumption. Similarly, $B$ is a tunable penalty weight that reduces the score in proportion to the packet loss rate, ensuring stability when loss is high.
III-C Offline Network Simulator
To employ a PPO agent as the optimizer, we must first train it. Training such an agent typically requires thousands of episodes, each consisting of multiple steps. This process becomes especially challenging when the reward for an action cannot be obtained immediately. In our case, at least three seconds of transfer data must be observed to compute the parameters of the utility function for a single step. As a result, one episode may take 30 seconds to one minute. Even under optimistic assumptions, no more than 3,000 episodes could be completed in a day—insufficient for training a PPO agent effectively. Moreover, in many practical scenarios, intended transfers may last less than one hour, making it impractical to dedicate a full day solely for training.
To address this, we developed an offline simulator capable of reproducing the necessary dynamics in less than four minutes.
The simulator, described in Algorithm LABEL:algo:simulator, models all essential aspects of data transfer in a simplified environment. It assumes an infinite pool of files. The files are no larger than the predefined chunk size to ensure parallelism is respected. The simulation is initialized with four parameters: network bandwidth, maximum throughput per thread, initial concurrency, and simulation duration. These values define the environment in which the agent interacts. Whenever the agent predicts a new action, the simulator triggers a call to the GET_UTILITYGET\_UTILITY method. Instead of instantiating actual concurrent transfers, the simulator emulates them using queues. Initially, the queue is populated with the current concurrency level. Each element of the queue, representing one concurrency thread, contains the current time and the remaining bandwidth available for that thread during a one-second interval. Processing proceeds by popping elements from the queue and invoking the TASK method with the time and per-thread bandwidth. This method emulates the time required to send a chunk and returns the next time the thread will begin sending as well as its updated remaining bandwidth. If the next start time exceeds the probing window, the thread is reinserted into the queue. After the queue is fully processed, the throughput is normalized to compute the reward, and the internal state of the simulator is updated accordingly.
III-D PPO Agent Architecture
Policy-driven deep reinforcement learning (DRL) algorithms are well-suited for problems where the environment changes rapidly and a generalizable actor is required. Among these, we adopt Proximal Policy Optimization (PPO) [schulman2017ppo], one of the most widely used policy-based reinforcement learning algorithms due to its balance of stability and sample efficiency. Our PPO agent follows the standard PPO architecture, with the key components highlighted in this subsection.
III-D1 State Space
Designing the state space is a crucial step, as it determines how the agent perceives the environment. If irrelevant components are included, the agent may learn misleading correlations with the utility function. At the same time, the state must be expressive enough so that distinct real-world scenarios do not collapse into the same representation. For instance, using only throughput as the state would be problematic, since the same throughput can occur under multiple concurrency settings that require different actions. Considering these factors, we configure the state with four variables. They are throughput, concurrency, throughput difference from the previous step, and concurrency difference from the previous step. Together, these features define a state representation that can be reliably mapped to an action without confusing the agent.
III-D2 Action Space
In our design, concurrency is selected as the action for the PPO agent. This choice allows the agent to directly map the observed state to the concurrency level, which is the single most important factor in determining the utility value. Other possible design choices do not directly map to the utility; for example, if we define concurrency change as the action, it only modifies concurrency, which in turn affects the utility. Learning this more complex relation requires additional time to converge.
III-D3 Policy Network
The policy network predicts actions directly from the current state using fully connected layers with residual connections. The input is first embedded into a 256-dimensional space through a linear layer with tanh activation, then passed through three residual blocks. Each block consists of two linear layers with normalization, ReLU, and a skip connection to aid gradient flow and capture complex state representations. The final output is processed by a tanh layer and a linear transformation to compute the mean of the action distribution. A trainable log–standard-deviation parameter, clamped and exponentiated, provides the variance, enabling sampling from a normal distribution to balance deterministic mapping with uncertainty.
III-D4 Value Network
The value network estimates the expected return for a state, essential for PPO advantage calculation. The state is projected into a 256-dimensional feature space using a linear layer with tanh just like the policy network, followed by two residual blocks with Tanh activations and skip connections to preserve gradient flow. The output is then passed through a linear layer to produce a scalar return estimate. This residual-based design improves stability and accuracy in dynamic environments.
Figure 2: LDM introduces offline training of a deep reinforcement learning agent to quickly learn the behavior of the real environment.
III-E The LDM Workflow
After describing the low-level components, we now present the high-level system design of LDM, illustrated in Figure 2. The workflow consists of three phases: the exploration phase, the offline training phase, and the production phase.
III-E1 Exploration and Logging
Before configuring the simulator, we run a short five-minute exploration phase. During this phase, memory-to-memory transfers are executed using random concurrency values. The throughput $T_{n}$ and concurrency $CC_{n}$ are logged by probing the network every second, where $n$ denotes the time in seconds.
From these measurements, we estimate the approximate network bandwidth $BW$ and per-concurrency throughput $TPT$ as
$BW=\max T_{n},\quad TPT=\max\frac{T_{n}}{CC_{n}},\quad$
To establish a convergence criterion for the simulator, we assume ideal linear scaling up to the available bandwidth with no packet loss. Under this assumption, the theoretical optimal concurrency $CC^{\star}$ and the maximum achievable reward $R_{\max}$ are defined as
$CC^{\star}=\frac{BW}{TPT},\quad R_{\max}=\frac{BW}{K^{CC^{\star}}}.$
III-E2 Training in the Simulated Environment
Once the simulator is configured with the data collected during the exploration phase, we begin training the PPO agent within this simulated environment.
At the start, the policy and value networks are initialized along with an experience buffer. In each episode, the simulator is reset and the agent interacts with the environment for a fixed number of steps. At each step, the policy outputs a Gaussian distribution over concurrency values, from which an action is sampled and executed. The resulting state, reward, and transition are recorded, and episode rewards are accumulated until termination. After each episode, discounted returns are computed and used to update the networks. The policy is optimized using the clipped surrogate objective to ensure stable learning, while the value network is trained with mean squared error. An entropy regularization term is included to encourage exploration. Both networks are jointly updated using the Adam optimizer, and the old policy is replaced with the updated one. Training proceeds until convergence is achieved, defined as reaching at least 90% of the theoretical maximum reward $R_{\max}$. To get the full potential of the agent we keep the training loop running till 1000 episodes pass without any improvement. The best-performing policy and value network are saved as the final model for adaptive thread allocation during transfers.
III-E3 Adaptive Concurrency in Production
In the production phase, we load the best-performing model obtained during training and begin real data transfers. At each step (rather than per episode), the policy network outputs a mean vector $\mu$ and log–standard deviation $\sigma$. An action is sampled from the corresponding diagonal Gaussian distribution and rounded to produce the predicted concurrency $CC_{n+1}$.
$CC_{n+1}=\mathrm{round}(\mathcal{N}(\mu,\sigma)).$
This value is then clamped to $1,CC_{\max}]$, where $CC_{\max}$ is a predefined parameter. Unlike in the training phase, the predicted concurrency is applied directly to stall or reopen workers rather than being passed to the simulator. The network is then probed with the updated concurrency setting to obtain the utility value and the new state. This process repeats continuously until all files in the dataset have been transferred.
IV Evaluation
(a) Concurrency Comparison
(b) Throughput Comparison
Figure 3: Performance comparison of LDM, Falcon, DRL, and GD with Heuristic with 4096 files sized 256 MB in Fabric-testbed. Only concurrency methods like DRL can reach the top speed, but LDMuses 25% less concurrency.
(a) Concurrency Comparison
(b) Throughput Comparison
Figure 4: The tail performance of the methods without parallelism (Falcon and DRL) shows the effect of having 4 large files of 55 GB with 804 files of 1 GB. They cannot utilize the available bandwidth.
In this section, we first discuss how offline training significantly accelerates the agent training process compared to its predecessor, DRL. We then demonstrate the impact of pipelining and parallelism, alongside concurrency, in improving data transfer performance.
IV-A Training in the Simulated Environment
After the logging and exploration phase, we proceed to the offline training of the agent within a simulated environment. We conducted experiments under different scenarios to estimate the time required to satisfy the convergence criteria described in Section III. Across 15 experiments, the mean number of episodes to convergence was 8,525 with a standard deviation of 5,065. In terms of wall-clock time, convergence required an average of 156 seconds with a standard deviation of 88 seconds. The maximum observed time was 222 seconds, while the minimum was 33 seconds.
[…]
TABLE I: Dataset Description
Dataset Total Size File Size Distribution
A (Uniform) 1 TB
4096 $\times$ 256 MB
B (Uniform and Large) 1 TB
804 $\times$ 1 GB
4 $\times$ 55 GB
C (Small) 128 GB
238,770 $\times$ {100 KB – 1 MB}
D (Mixed) 256 GB
2 $\times$ 12 GB
3872 $\times$ {10 MB – 100 MB}
23,624 $\times$ {100 KB – 1 MB}
IV-B Experimental Results
We evaluated LDM’s performance in terms of throughput maximization, early convergence, and stability across multiple emulated dataset scenarios. For benchmarking, we compared our approach against Falcon [arifuzzaman2023falcontpds], a recent state-of-the-art solution; Globus [globusonline], the most widely used transfer service; the DRL-based method by Hasibul et al. [drl]; and a gradient descent optimization strategy with heuristics enabled. In this section, we will refer to Hasibul’s method as DRL and the gradient descent optimization strategy with heuristics enabled as GD (Heuristic). Among the evaluated methods, only Globus operates in a static manner, whereas all others are adaptive. To evaluate the datasets using Globus, we initially attempted benchmarking with Globus Online. However, even after disabling verification, the transfer speed remained very slow. Therefore, for our experiments, we adopted the globus-url-copy utility from the community-maintained open-source Grid Community Toolkit (GCT 6.2). As discussed in Section II, we did not include FDT or MDTM in our comparisons due to the unavailability of their software.
We demonstrated the results through experiments conducted on Fabric [fabric-2019]. To emulate production-grade conditions, we utilized machines from two geographically distinct sites: one located in New York, USA (NEWY) and the other at CERN, Switzerland. Both machines were provisioned with 8 CPU cores, 64 GB of RAM, a Dell Express Flash P4510 1 TB SFF SSD, and an NVIDIA Mellanox ConnectX-6 NIC. While the physical port supports a line rate of 100 Gbps, in practice, we observed a maximum transfer speed of 58 Gbps.
[…]
TABLE II: End-to-End Transfer Speed Comparison Across Datasets (Gbps)
Model Dataset A Dataset B Dataset C Dataset D
Throughput Concurrency Throughput Concurrency Throughput Concurrency Throughput Concurrency
Globus 4.1 $\pm$ 0.1 32.0 $\pm$ 0.0 5.3 $\pm$ 0.2 32.0 $\pm$ 0.0 2.5 $\pm$ 0.1 64.0 $\pm$ 0.0 5.2 $\pm$ 0.6 64.0 $\pm$ 0.0
Falcon 23.7 $\pm$ 1.4 16.4 $\pm$ 6.0 22.2 $\pm$ 1.4 11.8 $\pm$ 3.5 0.7 $\pm$ 0.1 27.6 $\pm$ 13.1 3.4 $\pm$ 0.5 14.4 $\pm$ 5.9
DRL 39.5 $\pm$ 0.1 20.7 $\pm$ 2.4 24.1 $\pm$ 7.0 16.8 $\pm$ 3.3 1.2 $\pm$ 0.5 47.5 $\pm$ 2.9 4.0 $\pm$ 0.1 16.8 $\pm$ 1.5
LDM-GD 34.9 $\pm$ 1.7 11.7 $\pm$ 3.4 36.4 $\pm$ 3.0 12.3 $\pm$ 4.0 11.3 $\pm$ 0.6 2.4 $\pm$ 1.1 32.7 $\pm$ 2.5 3.8 $\pm$ 2.0
LDM 40.5 $\pm$ 1.2 15.1 $\pm$ 1.3 41.8 $\pm$ 0.2 15.5 $\pm$ 1.3 11.5 $\pm$ 0.3 3.8 $\pm$ 0.7 35.3 $\pm$ 2.7 8.6 $\pm$ 0.8
IV-B1 Effectiveness of Agent-Based Optimization
Dataset A consists of 4,096 files, each 256 MB in size. This dataset is intentionally designed so that the absence of pipelining and parallelism can be compensated with higher concurrency. This is possible because the number of files is not large enough for pipelining to become a significant bottleneck, and the file sizes are not big enough for parallelism to dominate performance.
To highlight the effectiveness of agent-based methods, we limited the per-concurrency network bandwidth to 3 Gbps. This setup allows us to observe how early convergence and stability influence throughput. The results are summarized in Table II, which reports the mean values along with their corresponding standard deviations.
Globus, by design, is static and operates differently from the other methods considered here. In our approach, parallelism is realized by fixing the chunk size, whereas Globus achieves parallelism by fixing the number of streams per file. In this experiment, Globus was configured with 32 concurrency and 4 parallel streams, resulting in a total of 128 TCP streams. Such a large number of streams, combined with file splitting and merging, can degrade performance. Hence, Globus is expected to underperform in all cases due to its design.
Among adaptive methods, DRL achieves nearly the same throughput as LDM, but it does so using 33% more concurrency, thereby validating our initial hypothesis. Falcon and GD (Heuristic) perform worse because online optimization requires additional time to identify the concurrency value that maximizes throughput. As demonstrated in Figure 3, Falcon and GD (Heuristic) only reached their mean throughput after 46 seconds and 40 seconds, respectively, compared to just 3 seconds and 6 seconds for DRL and LDM. This figure illustrates the best runs of each method in terms of throughput.
Finally, the standard deviation of concurrency further illustrates the effectiveness of agent-based methods. The standard deviations observed for PPO-based methods (DRL and LDM) are less than half of those for their gradient descent counterparts, demonstrating improved stability in optimization.
IV-B2 Handling Large Files with Parallelism
We designed Dataset B to include files that would take substantial time to transfer if not split into chunks. The per-concurrency bandwidth was throttled to 3 Gbps. Specifically, the dataset consists of four 55 GB files along with 804 files of size 1 GB. If the large files were to start simultaneously at the beginning, transferring a single 55 GB file with one thread would take approximately 55 GB * 8 / 3 Gbps $\approx$ 147 seconds even under maximum utilization of that thread. To highlight the necessity of parallelism, we deliberately structured the transfer queue such that the large files were not initiated at the same time. Methods with pipelining can efficiently handle these cases by chunking the large files and distributing them across multiple threads. This phenomenon is illustrated in Figure 4. The figure shows a steep drop in throughput for Falcon and DRL, both of which lack parallelism. Their throughput curves also exhibit a staircase-like pattern at the tail end. In DRL, one large file was completed before the drop, resulting in steps at 9, 6, and 3 Gbps. In Falcon, the steps occurred at 12 and 6 Gbps, indicating that two large files were completed in close succession.
This bottleneck directly impacts overall throughput, as reported in Table II. In this scenario, LDMclearly outperforms all baselines, achieving approximately 73% higher throughput than the state-of-the-art DRL.
IV-B3 Effect of Pipelining on Small Files
Dataset C is designed to demonstrate the effect of pipelining. This dataset consists of 238,770 files with sizes ranging from 100 KB to 1 MB. In the earlier cases, the lack of pipelining did not significantly degrade performance because the number of socket closures and reconnections was relatively small. However, with approximately fifty times more files than Dataset A, the overhead of repeatedly closing and reopening sockets becomes the dominant bottleneck. This effect is evident in the performance comparison shown in Table II. Interestingly, Globus outperforms Falcon and DRL in this scenario. Its fixed concurrency setting provides an advantage, since dynamically adjusting concurrency introduces additional socket operations. In contrast, methods that integrate pipelining are able to overcome this bottleneck and clearly outperform the others. Nevertheless, the performance of GD and LDM is somewhat slower than in other datasets, primarily due to the repeated encoding and decoding of file headers during transfers.
IV-B4 Mixed File Transfers Without Bandwidth Throttling
[…]
V Conclusion
Maximizing utilization of network bandwidth in high-speed networks is a critical challenge in the era of big data. Prior work has primarily addressed this problem using a single parameter, concurrency. In our study, we demonstrate that while concurrency alone can perform adequately in certain cases, it fails to generalize across diverse datasets. To address this limitation, we introduce two additional parameters—parallelism and pipelining—that consistently outperform existing methods, achieving up to $9.5\times$ Hotelsspeedup over state-of-the-art solutions. Furthermore, to reduce training overhead, we develop an offline network simulator that accelerates agent training by $2750\times$ compared to an online setting. Looking ahead, our goal is to design a generalizable, plug-and-play framework that eliminates the need for retraining whenever the underlying network environment changes.
Acknowledgement
The work in this study was supported by the NSF grant 2451376.